Digitel Electronics’ engineering and marketing departments have prepared forecasts for the development costs and operating profits of the next generation of their digital electrical meters. Development costs for each of the next three years will be $50,000.Manufacturing equipment costing $100,000 will be purchased near the end of Year 3. Annual profits for the normal five-year product life (Years 4 to 8 inclusive) are projected to be $80,000. The salvage value of the manufacturing equipment at the end of Year 8 is $20,000. Use NPV as your selection criteria. Should Digitel proceed with the product development if its annually compounded cost of capital is: a. 14%? b. 17%?
Answer #1

To decide whether the project is a good investment to undertake or not, we need to calculate the net present value (NPV) of project.

NPV is the sum of present value of all cash flows

If NPV is greater than zero then company should undertake this project as it is creating value for the company

To calculate the present value (PV) of a cash flow, we need to discount it by the required return.

PV = CFn/(1+r)n,

where CFn is cash flow in period n, r is required rate of return, n is time period

NPV = CF0 + CF1/(1+r)1 + CF2/(1+r)2 + ................... + CFn-1/(1+r)n-1 + CFn/(1+r)n

r is cost of capital in this case, as company will require to return at least cost of capital to generate value form this project.

Development cost in each of year 1, 2 and 3 = $50,000

Cost of manufacturing equipment = $100,000

Annual profits from year 4 to 8 = $80,000

Salvage value of manufacturing equipment = $20,000

So, cash flow in year 1 i.e. CF1 = $-50,000, negative sign indicates cash outflow

CF2 = $-50,000

CF3 = -50,000-100,000 = $-150,000

CF4 = $80,000

CF5 = $80,000

CF6 = $80,000

CF7 = $80,000

CF8 = $80,000+20,000 = $100,000

a) Cost of capital = 14%

Now putting all the values in NPV equation

NPV = -50,000/(1+0.14)1 -  50,000/(1+0.14)2 - 150,000/(1+0.14)3 + 80,000/(1+0.14)4 + 80,000/(1+0.14)5 + 80,000/(1+0.14)6 + 80,000/(1+0.14)7 + 100,000/(1+0.14)8

Solving this we get NPV as $8,810.98

As NPV > 0, therefore we should proceed with the development

b) Cost of capital = 17%

Now putting all the values in NPV equation

NPV = -50,000/(1+0.17)1 -  50,000/(1+0.17)2 - 150,000/(1+0.17)3 + 80,000/(1+0.17)4 + 80,000/(1+0.17)5 + 80,000/(1+0.17)6 + 80,000/(1+0.17)7 + 100,000/(1+0.17)8

Solving this we get NPV as $-7,414.45

As NPV < 0, therefore we should not proceed with the development